# A scalable quantum-neural hybrid variational algorithm for ground state estimation

### v1

@@ Introduction
Variational quantum eigensolver (VQE) [1, 2] is a key development in the field of quantum algorithms. VQE addresses one of the most challenging problems in quantum chemistry—calculating the ground state energy of molecular systems [3, 4, 5]. It employs a hybrid approach, using quantum computers to provide a parametrized quantum circuit (PQC) [6, 7] for evaluating expectation values, while classical computers optimize the circuit parameters. Significance of VQE lies in its scalability and adaptability to noisy intermediate-scale quantum (NISQ) hardwares [8], making it more hardware-friendly, especially in terms of circuit depths, compared to other ground-state estimating algorithms based on quantum phase estimation [9, 10].
@@
Although their expressiveness is limited, it has been shown that classical post-processing can enhance their capabilities [13, 17, 18, 19]. A novel approach, the variational quantum-neural hybrid eigensolver (VQNHE) [13] exploits neural network to apply additional transformations to the quantum state. The neural network processes each quantum circuit measurement as a binary string, producing a single-numbered output without exponential computational overhead. The expectation value of the Hamiltonian is then calculated by combining measurement results with neural network outputs. This method, using a statevector simulator, has demonstrated significant improvements for transverse-field Ising model (TFIM) [20] and molecular Hamiltonian simulations compared to original VQE using hardware-efficient ansatze.
Despite improved performance achieved with VQNHE, this combination of quantum and classical processes has a critical defect. We have discovered that it can optimize the neural network to yield a very large negative expectation value of the given Hamiltonian regardless of the actual ground state energy, and that such phenomenon can only be prevented with exponential amount of quantum circuit measurements, causing scalability bottlenecks in terms of practical implementation. To address the challenge at hand, we propose a novel way of utilizing neural network to perform a nontrivial unitary transformation in a complex domain. This structure circumvents the need for normalization, which has been the primary reason for the critical divergence during neural network training. With this advanced method, we eliminate the need for an exponentially large number of circuit shots, thereby improving its computational efficiency.
@@
The VQNHE algorithm appends neural network to VQE by feeding the quantum circuits output in binary bit strings to the neural network, as shown in Fig. 1. The network maps an $n$-bit binary string to a floating point number to yield the following transformation:
@@
A method to efficiently evaluate the Pauli terms has also been provided. The qubit corresponding to the first X or Y term in the Pauli term is designated as the star qubit $q^{*}$. Here on, define $s^{\prime}\in\{0,1\}^{\otimes n}$ as the bit string with its star qubit set to 00 while leaving the other qubits, i.e. $s=q_{0}q_{1}...q^{*}...q_{n}\rightarrow s^{\prime}=q_{0}q_{1}...0...q_{n}$. Also define the tilde transformation $s\rightarrow\tilde{s}_{P}$ as follows: for a Pauli letter in the Pauli string $P\in\{I,X,Y,Z\}$, flip the bit if it is either $X$ or $Y$, and leave otherwise. For example, for a bit string $s=110$ and a Pauli string $P=IXY$, $\tilde{s}_{P}=101$. As the Pauli letter corresponding to the star qubit is either $X$ or $Y$, this transformation always flips the star qubit $q^{*}$. Since in $s^{\prime}$ the star qubit value is 00 at all times, the star qubit value in $\tilde{s^{\prime}}_{P}$ is always $1$. The expectation value for each Pauli term $\hat{P}$ becomes
@@
where $|\psi_{m}(P)\rangle$ represents the ansatz with an additional measurement circuit determined by the Pauli string $P$. The measurement circuit applies controlled-$X$/$Y$ gates on the non-star qubits corresponding to the $X$ or $Y$ strings with the star qubit as the control, followed by basis transformations for measurement on the given Pauli basis. Such transformation yields diagonalization of the Pauli string [13], allowing for efficient evaluation of the expectation value with respect to the transformed state. The overall workflow is graphically described in Fig. 1. In the following sections, we call the ansatz circuits with the attached transformations as measurement circuits.
@@
The neural transformation of the state generated from the ansatz is in general a non-unitary transformation, regardless of the output range of the neural network. To take care of this, VQNHE have included normalization in the final step of evaluating the ground state energy in Eq. (3). However, we have discovered that successful normalization requires an extremely large number of shots, growing exponentially with the number of qubits (See Appendix). This issue arises because the loss function involves bit strings from both the ansatz and measurement circuits, and successful normalization requires that all contributing bit strings be sampled at least once. If any bit string appearing in the numerator of the loss function, originating from the measurement-combined circuits, is absent from the ansatz measurement outcomes, the neural network can exploit the mismatch and induce divergence in the loss. Using the coupon collector analysis, ensuring full coverage demands $O(2^{n}\log N_{M})$ shots on the ansatz circuit, $N_{M}$ representing the set of bit strings that serve as inputs to the neural network in the numerator.
Figure 2: VQNHE implementation of a 7-site TFIM with 7 qubits. (a) Training of the neural network in VQNHE. The vertical axis shows the loss function—the expectation value of the Hamiltonian—on a log scale. The quantum circuit simulation uses the Qiskit sampler with 500 shots per circuit. The inset highlights the region between the lowest bare VQE value and the exact ground state energy; values below this are invalid. (b) Output of the trained neural network after 200 epochs. The horizontal axis shows bit strings (in decimal), and the vertical axis shows neural network outputs. Blue lines correspond to measured bit strings, red lines to unmeasured ones. Dots mark the network outputs. While most values lie near $10^{-6}$, extreme values (above $10^{-2}$, shown in red) appear. Since nearly all bit strings are covered by the measurement circuits, the red dots represent can be thought only to have appeared in the numerator.
@@
The behavior is clearly demonstrated in Fig. 3. Note that quantum circuit measurements without all $322^{5}=$ 32 bit strings present were excluded from the sample, as they experience severe divergence as mentioned in the previous section. As the overall variance scales with $\frac{1}{N}$ and $\frac{1}{N^{2}}$ terms, the VQNHE results converge to the exact VQNHE value (shown in blue line) at larger $N$. However, without sufficient number of shots, it falls below the exact ground state energy, which is not a valid solution. One thing to notice is the existence of poorly trained results lying significantly above the exact VQE results due to the neural network stuck in a wrong minimum due to discrepancies of the measured probabilities and the actual ones. In overall, even with all bit strings measured at least once, which already costs exponential amount of measurements, the incongruity between the shot-based VQNHE and its exact version is still large, resulting in even worse approximation to the ground state energy than the VQE alone at times.
@@
Fig. 4 compares U-VQNHE with the standard VQNHE, both with statevector simulator and shot-based sampler for the execution of quantum circuits. The dashed lines show the results from the statevector simulator [24]. Sampler results with finite shots are shown in solid lines. As previously shown, shot-based VQNHE quickly plunges to large negative values, whereas U-VQNHE stays in the region between the exact ground state energy and the optimized value of VQE even with the sampler. Although U-VQNHE does not achieve expectation values as low as those from VQNHE, it is the only variant suitable for implementation on shot-based quantum hardware.
In addition to relaxing the requirement that all $2^{n}$ basis bit strings of $\mathcal{H}_{n}$ be present, thereby avoiding the divergence issues seen in VQNHE, U-VQNHE also exhibits smaller deviations from the exact ground state energy when all bit strings are included. Fig. 4(c) shows how the number of shots affects each algorithm. The key factor making U-VQNHE more stable is that the variance of VQNHE shows dependencies on magnitudes of the neural network outputs $f(s)$. Standard VQNHE has no specific restrictions on their values, whereas U-VQNHE restricts their absolute values to $1$. When scaling up the system, the uncertainty due to lack of shots is a huge bottleneck, and U-VQNHE mitigates the uncertainty of VQNHE.
@@
In this research, it has been shown that while VQNHE possesses a powerful ability to enhance the expressiveness of VQEs with hardware-efficient ansatze, actual implementation of the algorithm requires exponential amount of measurements on the quantum circuits in order for the optimization to stay within the bound. Moreover, even if all $2^{n}$ bit strings expressed with $n$ qubits are present out of the ansatz circuit, which is a condition that must be strictly kept to prevent the divergence, VQNHE failed to produce accurate values for the ground state of $n$-site TFIM due to the inherently large variance. It comes from the fact that the values of the neural network can be large.
To address these issues to apply the neural network in a fully scalable fashion, we have suggested unitary-VQNHE. The main focus is to avoid the necessity to use the normalization at all, which have been the cause of strange behaviors, while achieving comparable expressiveness. To do so, we have redesigned the neural network to yield a complex value. The suggested algorithm has been shown to yield the ground state energy of 5- and 12-site TFIM stably and limit the variance owing to the unitarity of the neural network outputs. By resolving scalability issues of VQNHE while maintaining its advantages in resource efficiency, U-VQNHE not only enhances stability and applicability of the direct applications of VQNHE [26, 27], but also possesses potential for quantum advantage in the near future for ground state estimation of complex Hamiltonian with simple ansatz structures.
@@ Appendix D Divergence of VQNHE with positive neural network values
At the baseline, there must exist an element in $B_{M}\backslash B_{a}$, which we denote as $s^{*}$. This is the divergence condition for general neural network. Now, viewing in terms of $f(s^{*})$, in the Loss function each of it is multiplied by
@@
where in this case, the star qubit can either take 00 or $1$. For elements in $B_{M}\backslash B_{a}$, the condition for divergence under positive-valued neural network is the following: given a Hamiltonian and probability distribution, there exists at least one element $s^{**}$ such that $h_{P}(-1)^{q^{*}}f(\tilde{s^{**}_{P}})p_{m}(s;P)<0$ for some neural network $f(s)$. In other words, if the neural network can manage to find a way to yield $f(\tilde{s^{**}_{P})}$ values such that this value is negative, positively growing $f(s^{**})$ results in negative increase of the loss function, causing the divergence.
@@ Appendix E Derivation of the variance of VQNHE
Evaluating Eq. (23) and Eq. (26) for the Pauli terms composing the Hamlitonian, one can evaluate the approximate variance of $\langle H\rangle-\langle H\rangle_{m}$.
@@
Note that this is a slight overestimation of the variance, since $\mathbb{E}[1/X]>1/\mathbb{E}[X]$. As the entire variance is mainly proportional to $1/N$ with additional $1/N^{2}$ terms due to the variance of the denominator, the entire variance disappears at infinite shots, as expected. In addition, in overall, all terms show heavy dependence on the values of $f(s)$. Similar to what has been asserted in the authors of Ref. [13], wide distribution of the values $f(s)$ is a critical reason for large deviation of VQNHE. On the other hand, although the exact formula does not directly apply, U-VQNHE heavily constrains the values of output of the neural network, reducing the deviation in general.
@@
VQNHE is not the only algorithm that utilizes non-unitary transformation onto VQE. There have been several researches that attempts to apply additional gates and post-processing in a non-unitary fashion in order to achieve better performance in terms of accuracy, especially for the eigensolvers with hardware-efficient ansatze. We briefly review potentials for measurement-derived divergence discussed in our work for several researches that utilize non-unitary transformations onto the states and discuss their resource efficiencies compared to our work.
@@
Going through the scenarios of divergence in the paper, the first obvious possibility is that viewing the parameters $\vec{\alpha},\vec{\lambda}$ as variables, any coefficient on the denominator yields zero and the corresponding one on the numerator gives nonzero value. This corresponds to the case where bit strings out of measurement process are missing on the denominator side, causing severe failure of the algorithm. Nevertheless, this is not a realistic situation in terms of nu-VQE. In order to satisfy the conditions, it must be the case where $\exists i$ such that $\langle\psi|\hat{Z_{i}}|\psi\rangle=0,\langle\psi|\hat{P}\hat{Z_{i}}|\psi\rangle\neq 0$ or $\exists i,j$ such that $\langle\psi|\hat{Z_{i}}\hat{Z_{j}}|\psi\rangle=0,\langle\psi|\hat{P}\hat{Z_{i}}\hat{Z_{j}}|\psi\rangle\neq 0$. Although not theoretically impossible, getting an exact zero as an expectation value is highly unlikely, unless one happens to prepare an exactly balanced state as $|\psi\rangle$ or the numbers of bit strings that yield $-1$ and $+1$ for $Z_{i}$ are the same out of measurements. In such rare cases, it might experience similar failure to VQNHE, although it does not count as a major flaw of the algorithm.
Despite its idea of applying a non-unitary transformation without significant risk of divergence and thus without concerns of scalability bottleneck due to number of shots, nu-VQE requires much more of quatnum circuit evaluations compared to VQNHE or U-VQNHE. As written in Eq. (30), in order to provide the expectation values to all the terms, it requires at most $O(n^{4})$ quantum circuit evaluations for each Pauli term comprising the target Hamiltonian. Note that this is minimal overhead of using the Jastrow factor at its linear approximation. Any higher-ordered terms require much more circuits, and the exact Jastrow factor demands exponential of them. Thus, VQNHE and U-VQNHE require much less computational resources in terms of the quantum computations, while requiring only a polynomial scale of classical resources.
@@
Note that it captures the second-order terms of Eq. (29) while not exploiting its linear approximation form. In Ref. [18] it is argued that the nonunitary Jastrow operator effectively filter out components that yield high energy eigenvalues.
Although the detailed implementation and approximation details differ, JQC and nu-VQE actually end up with the same procedure from the expectation value point-of-view. The difference comes from the target on which the Jastrow operator is acts. Thus, in terms of possible divergence, same arguments on nu-VQE applies for non-unitary transformations by projective mapping of states. In addition, in terms of resource efficiency, it experiences the same overhead in the number of quantum circuits required, making VQNHE and U-VQNHE superior considering the resource overhead.
@@
where $U$ is a unitary operator and $|0\rangle$ represents the vacuum state, written in the second quantization picture. It can be a fixed operator or be trained with simple ansatze, just as hardware-efficient ansatz block is trained in VQNHE before optimizing parameters of the neural network. The result of the expectation value of the given Hamiltonian can then be written as
@@
In short, there are several other variants of VQE that utilizes non-unitary transformation to enhance the expressiveness. While algorithms that makes use of Jastrow factor to modify the quantum state mapping[18] or the Hamiltonian[19] are unlikely to experience the divergence during the optimization of the parameters of the non-unitary transformations, they require much more quantum resources than VQNHE or U-VQNHE and thus are not as resource-efficient. On the other hand, CVQE[17] displays a lot of resemblance to VQNHE in terms of how they evaluate the expectation value with respect to the parameters, and it has been shown that CVQE also requires exponential amount of shots to be taken to evaluate the denominator to prevent itself from perilous divergence. Thus, no algorithm among those listed above presents better resource-efficiency than U-VQNHE considering both the number of quantum circuits to be evaluated and the number of shots to be taken.

### v2

@@ Introduction
Variational quantum eigensolver (VQE) [1, 2] is a key development in the field of quantum algorithms. VQE addresses one of the most challenging problems in quantum chemistry—calculating the ground state energy of molecular systems [3, 4, 5]. It employs a hybrid approach, using quantum computers to provide a parametrized quantum circuit (PQC) [6, 7] for evaluating expectation values, while classical computers optimize the circuit parameters. Significance of VQE lies in its scalability and adaptability to noisy intermediate-scale quantum (NISQ) hardware [8], making it more hardware-friendly, especially in terms of circuit depths, compared to other ground-state estimating algorithms based on quantum phase estimation [9, 10].
@@
Although their expressiveness is limited, it has been shown that classical post-processing can enhance their capabilities [13, 17, 18, 19]. A novel approach, the variational quantum-neural hybrid eigensolver (VQNHE) [13], exploits a neural network to apply additional transformations to the quantum state. The neural network processes each quantum circuit measurement as a binary string, producing a single-numbered output without exponential computational overhead. The expectation value of the Hamiltonian is then calculated by combining measurement results with neural network outputs. This method, using a statevector simulator, has demonstrated significant improvements for transverse-field Ising model (TFIM) [20] and molecular Hamiltonian simulations compared to the original VQE using hardware-efficient ansatze.
Despite improved performance achieved with VQNHE, this combination of quantum and classical processes has a critical defect. We have discovered that it can optimize the neural network to yield a very large negative expectation value of the given Hamiltonian regardless of the actual ground state energy, and that such a phenomenon can only be prevented with an exponential number of quantum circuit measurements, causing scalability bottlenecks in terms of practical implementation. To address the challenge at hand, we propose a novel way of utilizing a neural network to perform a nontrivial unitary transformation in a complex domain. This structure circumvents the need for normalization, which has been the primary reason for the critical divergence during neural network training. With this advanced method, we eliminate the need for an exponentially large number of circuit shots, thereby improving its computational efficiency.
@@
The VQNHE algorithm appends a neural network to VQE by feeding the quantum circuits output in binary bit strings to the neural network, as shown in Fig. 1. The network maps an $n$-bit binary string to a floating point number to yield the following transformation:
@@
A method to efficiently evaluate the Pauli terms has also been provided. The qubit corresponding to the first X or Y term in the Pauli term is designated as the star qubit $q^{*}$. Here on, define $s^{\prime}\in\{0,1\}^{\otimes n}$ as the bit string with its star qubit set to 0 while leaving the other qubits, i.e. $s=q_{0}q_{1}...q^{*}...q_{n}\rightarrow s^{\prime}=q_{0}q_{1}...0...q_{n}$. Also define the tilde transformation $s\rightarrow\tilde{s}_{P}$ as follows: for a Pauli letter in the Pauli string $P\in\{I,X,Y,Z\}$, flip the bit if it is either $X$ or $Y$, and leave otherwise. For example, for a bit string $s=110$ and a Pauli string $P=IXY$, $\tilde{s}_{P}=101$. As the Pauli letter corresponding to the star qubit is either $X$ or $Y$, this transformation always flips the star qubit $q^{*}$. Since in $s^{\prime}$ the star qubit value is 0 at all times, the star qubit value in $\tilde{s^{\prime}}_{P}$ is always 111. The expectation value for each Pauli term $\hat{P}$ becomes
@@
[…]
@@
The neural transformation of the state generated from the ansatz is in general a non-unitary transformation, regardless of the output range of the neural network. To take care of this, VQNHE has included normalization in the final step of evaluating the ground state energy in Eq. (3). However, we have discovered that successful normalization requires an extremely large number of shots, growing exponentially with the number of qubits (See Appendix). This issue arises because the loss function involves bit strings from both the ansatz and measurement circuits, and successful normalization requires that all contributing bit strings be sampled at least once. If any bit string appearing in the numerator of the loss function, originating from the measurement-combined circuits, is absent from the ansatz measurement outcomes, the neural network can exploit the mismatch and induce divergence in the loss. Using the coupon collector analysis, ensuring full coverage demands $O(2^{n}\log N_{M})$ shots on the ansatz circuit, $N_{M}$ representing the set of bit strings that serve as inputs to the neural network in the numerator.
Figure 2: VQNHE implementation of a 7-site TFIM with 7 qubits. (a) Training of the neural network in VQNHE. The vertical axis shows the loss function—the expectation value of the Hamiltonian—on a log scale. The quantum circuit simulation uses the Qiskit sampler with 500 shots per circuit. The inset highlights the region between the lowest energy from the bare VQE (i.e., VQE without a neural network) and the exact ground state energy; values below this are invalid. (b) Output of the trained neural network after 200 epochs. The horizontal axis shows bit strings (in decimal), and the vertical axis shows neural network outputs. Blue lines mark bit strings that were sampled (observed) from the ansatz, and red lines mark bit strings that were never observed. Dots mark the network output values. While most values lie near $10^{-6}$ (shown by black dots), extreme values (above $10^{-2}$, shown in red) appear. The neural network assigns extremely large output values to some of those unobserved (red) strings, effectively contributing only to the numerator of the loss (since they are absent in the denominator).
@@
The behavior is clearly demonstrated in Fig. 3. Note that quantum circuit measurements without all $2^{5}=32$ bit strings present were excluded from the sample, as they experience severe divergence as mentioned in the previous section. As the overall variance scales with $\frac{1}{N}$ and $\frac{1}{N^{2}}$ terms, the VQNHE results converge to the exact VQNHE value (shown in blue line) at larger $N$. However, without sufficient number of shots, it falls below the exact ground state energy, which is not a valid solution. One thing to notice is the existence of poorly trained results lying significantly above the exact VQE results due to the neural network stuck in a wrong minimum due to discrepancies of the measured probabilities and the actual ones. Overall, even with all bit strings measured at least once, which already costs an exponentially large number of measurements, the incongruity between the shot-based VQNHE and its exact version is still large, resulting in even worse approximation to the ground state energy than the VQE alone at times.
@@
Fig. 4 compares U-VQNHE with the standard VQNHE, both with statevector simulator and shot-based sampler for the execution of quantum circuits. The dashed lines show the results from the statevector simulator [24]. Sampler results with finite shots are shown in solid lines. As previously shown, shot-based VQNHE quickly plunges to large negative values, whereas U-VQNHE stays in the region between the exact ground state energy and the optimized value of VQE even with the sampler. Although U-VQNHE does not achieve expectation values as low as those from VQNHE, it is far more suitable for implementation on shot-based quantum hardware given current hardware limitations.
In addition to relaxing the requirement that all $2^{n}$ basis bit strings of $\mathcal{H}_{n}$ be present, thereby avoiding the divergence issues seen in VQNHE, U-VQNHE also exhibits smaller deviations from the exact ground state energy when all bit strings are included. Fig. 4(c) shows how the number of shots affects each algorithm. The key factor making U-VQNHE more stable is that the variance of VQNHE shows dependencies on magnitudes of the neural network outputs $f(s)$. Standard VQNHE has no specific restrictions on their values, whereas U-VQNHE restricts their absolute values to 111. When scaling up the system, the uncertainty due to lack of shots is a huge bottleneck, and U-VQNHE mitigates the uncertainty of VQNHE.
@@
In this research, it has been shown that while VQNHE possesses a powerful ability to enhance the expressiveness of VQEs with hardware-efficient ansatze, actual implementation of the algorithm requires an exponentially large number of measurements on the quantum circuits in order for the optimization to stay within the bound. Moreover, even if all $2^{n}$ bit strings expressed with $n$ qubits are present out of the ansatz circuit, which is a condition that must be strictly kept to prevent the divergence, VQNHE failed to produce accurate values for the ground state of $n$-site TFIM due to the inherently large variance. It comes from the fact that the values of the neural network can be large.
To address these issues to apply the neural network in a fully scalable fashion, we have suggested unitary-VQNHE. The main focus is to avoid the necessity to use the normalization at all, which have been the cause of anomalous behavior, while achieving comparable expressiveness. To do so, we have redesigned the neural network to yield a complex value. The suggested algorithm has been shown to yield the ground state energy of 5- and 12-site TFIM stably and limit the variance owing to the unitarity of the neural network outputs. By resolving scalability issues of VQNHE while maintaining its advantages in resource efficiency, U-VQNHE not only enhances stability and applicability of the direct applications of VQNHE [26, 27], but also possesses potential for quantum advantage in the near future for ground state estimation of complex Hamiltonian with simple ansatz structures.
@@ Appendix D Divergence of VQNHE with positive neural network values
[…]
@@
[…]
@@ Appendix E Derivation of the variance of VQNHE
Evaluating Eq. (23) and Eq. (26) for the Pauli terms composing the Hamiltonian, one can evaluate the approximate variance of $\langle H\rangle-\langle H\rangle_{m}$.
@@
[…]
@@
VQNHE is not the only algorithm that utilizes non-unitary transformation onto VQE. Several research efforts attempt to apply additional gates and post-processing in a non-unitary fashion in order to achieve better performance in terms of accuracy, especially for the eigensolvers with hardware-efficient ansatze. We briefly review potentials for measurement-derived divergence discussed in our work for several stuides that utilize non-unitary transformations onto the states and discuss their resource efficiencies compared to our work.
@@
Going through the scenarios of divergence in the paper, the first obvious possibility is that viewing the parameters $\vec{\alpha},\vec{\lambda}$ as variables, any coefficient on the denominator yields zero and the corresponding one on the numerator gives nonzero value. This corresponds to the case where bit strings out of measurement process are missing on the denominator side, causing severe failure of the algorithm. Nevertheless, this is not a realistic situation in terms of nu-VQE. In order to satisfy the conditions, it must be the case where $\exists i$ such that $\langle\psi|\hat{Z_{i}}|\psi\rangle=0,\langle\psi|\hat{P}\hat{Z_{i}}|\psi\rangle\neq 0$ or $\exists i,j$ such that $\langle\psi|\hat{Z_{i}}\hat{Z_{j}}|\psi\rangle=0,\langle\psi|\hat{P}\hat{Z_{i}}\hat{Z_{j}}|\psi\rangle\neq 0$. Although not theoretically impossible, getting an exact zero as an expectation value is highly unlikely, unless one happens to prepare an exactly balanced state as $|\psi\rangle$ or the number of measurement outcomes yielding –1 for $Z_{i}$ equals the number yielding +1. In such rare cases, it might experience similar failure to VQNHE, although this is not considered a major flaw of the algorithm.
Despite its idea of applying a non-unitary transformation without significant risk of divergence and thus without concerns of scalability bottleneck due to number of shots, nu-VQE requires many more quantum circuit evaluations than VQNHE or U-VQNHE. As written in Eq. (30), in order to provide the expectation values to all the terms, it requires at most $O(n^{4})$ quantum circuit evaluations for each Pauli term comprising the target Hamiltonian. Note that this is minimal overhead of using the Jastrow factor at its linear approximation. Any higher-ordered terms require much more circuits, and the exact Jastrow factor demands exponential of them. Thus, VQNHE and U-VQNHE require much less computational resources in terms of the quantum computations, while requiring only polynomially scaling classical resources.
@@
Note that it captures the second-order terms of Eq. (29) while not exploiting its linear approximation form. In Ref. [18] it is argued that the nonunitary Jastrow operator effectively filters out components that yield high energy eigenvalues.
Although the detailed implementation and approximation details differ, JQC and nu-VQE actually end up with the same procedure from the expectation value point-of-view. The difference comes from the target on which the Jastrow operator acts. Thus, in terms of possible divergence, the same arguments on nu-VQE apply for non-unitary transformations by projective mapping of states. In addition, in terms of resource efficiency, it experiences the same overhead in the number of quantum circuits required, making VQNHE and U-VQNHE superior considering the resource overhead.
@@
where $U$ is a unitary operator and |0|0$\rangle$| 0 ⟩ represents the vacuum state, written in the second quantization picture. It can be a fixed operator or be trained with simple ansatze, just as hardware-efficient ansatz block is trained in VQNHE before optimizing parameters of the neural network. The result of the expectation value of the given Hamiltonian can then be written as
@@
In short, there are several other variants of VQE that utilizes non-unitary transformation to enhance the expressiveness. While algorithms that makes use of Jastrow factor to modify the quantum state mapping[18] or the Hamiltonian[19] are unlikely to experience the divergence during the optimization of the parameters of the non-unitary transformations, they require much more quantum resources than VQNHE or U-VQNHE and thus are not as resource-efficient. On the other hand, CVQE[17] displays a lot of resemblance to VQNHE in terms of how they evaluate the expectation value with respect to the parameters, and our analysis indicates that CVQE also requires an exponential amount of shots to be taken to evaluate the denominator to prevent itself from divergence issues. Thus, no algorithm among those listed above presents better resource-efficiency than U-VQNHE considering both the number of quantum circuits to be evaluated and the number of shots to be taken.